es of stock at a price of $35 and an initial margin of 70 percent. If the maintenance margin is 40 percent, at what share price will you receive a margin call? What calculations. Round your answers to 2 decimal places.) account calculations. Round your answers to 2 decimal plac esyour ccount equlity at this stock price? (Do not round intermediate Margin call price Account equity
Answer #1

Proceeds from short sale = 650*$35 = $22,750

Initial deposit = $22,750 * 70% = $15,925

Account value = $22,750 + 15,925 = $38,675

Margin call price = $38,675 / [650 + (0.40 × 650)] = $42.50

Account equity = $38,675 – (650 × $42.50) = $11,050